OpenStudy (anonymous):

24a − 22 = −4(1 − 6a)

OpenStudy (anonymous):

all you need to do it distribute the -4(1-6a). so then you will have 24a-22=-4+24a then take and combined like terms -22+4=-18. then take 24a-24a=0. so it will end up being -18=0... if that's what the directions say